OpenStudy (lgbasallote):

an RC circuit with a 1 ohm resistance and 0.000001 F capacitor is driven by a voltage E = sin 100t. If the initial voltage across the capacitor is zero, find the subsequent resistor and capacitor voltages, and the current

OpenStudy (kropot72):

\[E=\sin 100t=\sin \omega t\] The capacitive reactance in the series circuit = \[\frac{1}{\omega C}=10,000 \Omega\] Therefore the analysis can be simplified by assuming the resistor value is zero ohms, the reason being the resistor value is negligible compared with the reactance of the capacitor. At time t the charge on the capacitor plates is \[Q=CE=C \sin \omega t\] \[\frac{dQ}{dt}=\omega C \cos \omega t=\omega C \sin (\omega t+\frac{\pi}{2})=I\]

OpenStudy (lgbasallote):

thanks @kropot72 i'll try that :D

OpenStudy (kropot72):

You're welcome :)